OpenStudy (anonymous):

y-5=3(x-2) in standard form?

hero (hero):

1. Add 5 to both sides: y = 3(x - 2) + 5 2. Simplify the right side

OpenStudy (anonymous):

y=3x-6=5?

OpenStudy (anonymous):

y=3x-6+5 i mean?

OpenStudy (anonymous):

y=3x-1. now to get this in standard form you have to move 3x to the other side.

OpenStudy (anonymous):

3x+y=1

OpenStudy (anonymous):

?

hero (hero):

No, you made a mistake there bro

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

when you move something to one side what happens to the sign

OpenStudy (anonymous):

I dont know I'm confused?

OpenStudy (anonymous):

Can you help me step by step?

OpenStudy (anonymous):

when moving something from one side of the equation to the other side you have to change the sign

OpenStudy (anonymous):